OpenStudy (anonymous):

Can someone teach me how to find the length of the diagonal when the length and width of the RECTANGLE is given .. :3 please!!!

OpenStudy (anonymous):

example is the length is 12 and width is 9

OpenStudy (anonymous):

use Pythagoras theorem... where the formula is : \[x = \sqrt{w ^{2}+l ^{2}}\]

OpenStudy (anonymous):

tnx ,, what about this,, the length is \[3\] \[\sqrt{2}\] and the width is 2\[\sqrt{3}\]

OpenStudy (anonymous):

3square root of 2 and 2 square root of 3

OpenStudy (anonymous):

Then it's squareroot of (3^2)+(2^2)

OpenStudy (anonymous):

which is squareroot of 13 which equals approx. 3,6

OpenStudy (anonymous):

its actually